3. A candy bar producer is producing a new candy bar that has a normal distribution with a mean of 43.4 grams and a standard deviation of .15 grams

a. If every bar is labeled as 43 grams, what proportion of the bars contain less than the labeled amount?

b. If only 1.5% of the candy bars exceed weight w, what is the value of w?
Answer #1

a) Proportion of bars that contain less than labeled amount

= P(X < 43)

= P(z < (43 - 43.4)/0.15)

= P(z < -2.67)

= 0.0038

b) z score corresponding to 1.5% area to the right tail = 2.17

Hence,

w = 43.4 - 2.17*0.15 = 43.73 grams